OpenStudy (dancergirl45):

PLEASE WILL FAN AND MEDAL!!!!!!!!!!!!!!! Two sets of equatic expressions are shown below in various forms: Line 1: x2 − x − 6 (x + 2)(x − 3) (x + 0.5)2 − 6.25 Line 2: x2 − x − 2 (x + 1)(x − 2) (x − 0.5)2 − 2.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (dancergirl45):

@YoungStudier

OpenStudy (dancergirl45):

@owikbenstan

OpenStudy (owikbenstan):

This is basically asking you to put a quadratic equation into all of it's different forms, I'm sure you know how to factor the expressions but the final step would be completing the square. To do this, do the following Step 1: Complete the Square \[x^2-x+(\frac{-1}{2})^2=+6+(\frac{-1}{6})^2\] Note that I took the original 6 on the left and added it to both sides, this will allow me to complete the square on the left side, which has been done in this step, also note that when you complete the square you must add the completed square on BOTH sides so that you do not change the equation. Step 2: Write the equation as a perfect square. \[(x-(\frac{1}{2}))^2=+6+\frac{1}{4}\] Step 3: \[(x-(\frac{1}{2}))^2-6-\frac{1}{4}=0\] this is the same as \[(x-0.5)^2-6.25 = 0\]

OpenStudy (dancergirl45):

so it would be: B. line 2 only

OpenStudy (dancergirl45):

right?

OpenStudy (owikbenstan):

Yes.

OpenStudy (dancergirl45):

thank you so much!!

OpenStudy (owikbenstan):

No problem. :)